# Uncertainty-Guided Optimization on Large Language Model Search Trees

### v1

@@ 3.3 Posterior beliefs over the optimal values
Notice that whenever a new node $x_{i}$ is added to the search tree, the likelihood associated with the path from the root to that node $c_{\boldsymbol{x}_{0}\rightarrow x_{i}}$ is fully observed—its distribution is simply a Dirac delta. This means, the joint distribution becomes $p(c_{\boldsymbol{x}_{0}\to x_{i}},\Delta_{i})=\delta(c_{\boldsymbol{x}_{0}\to x_{i}})\,p(\Delta_{i})$. Therefore, to sample $v_{x_{i}}$ given that we have observed $c_{\boldsymbol{x}_{0}\to x_{i}}$—i.e., sampling from the posterior $p(v_{x_{i}}\mid c_{\boldsymbol{x}_{0}\rightarrow x_{i}})$—it is sufficient to sample from $p(\Delta_{i})$ and then simply scale all samples by $c_{\boldsymbol{x}_{0}\rightarrow x_{i}}$. We leverage these posterior samples to make decision in exploring the tree (i.e., in expanding the nodes).
@@ 3.4 Decision making
while ^$\hat{\mathbb{P}}(c^{*}<v_{x_{0}^{*}})>\varepsilon$ do
@@
^$\hat{\mathbb{P}}(c^{*}<v_{\boldsymbol{x}_{0}^{*}})\leftarrow\frac{1}{N}\sum_{n=1}^{N}\mathbb{I}\left[c^{*}\leq(v_{\boldsymbol{x}_{0}^{*}})_{n}\right]$
@@
At each iteration, we use the posterior samples by following the steps of Monte-Carlo tree search (MCTS), without the expensive rollout step. Alg. 2 contains the pseudocode for the decoding portion of ULTS.
@@
Selection Starting from the root node $\boldsymbol{x}_{0}$, we recursively pick a child that has the highest empirical probability to be the best child, as computed based on the posterior samples. I.e. at a node $x_{i}$, we recursively maximize the following acquisition function $a_{i}(x_{c})$ over the children $x_{c}\in\mathrm{children}{(}x_{i})$:
@@
$a_{i}(x_{c})=\frac{1}{N}\sum_{n=1}^{N}\mathbb{I}\bigl{[}(v_{x_{c}^{*}})_{n}=\max_{x_{j}\in\mathrm{children}(x_{i})}(v_{x_{j}^{*}})_{n}\bigr{]},$ (1)
Intuitively, this means that we are estimating the probability that a child $x_{c}$ is the best child among other children of $x_{i}$. This recursive process is done until an unexpanded node, i.e. one that is in the current tree boundary $\mathcal{L}$, is selected.
@@
Different strategies can be used. The most straightforward is to use the children’s posterior samples themselves. This corresponds to simply setting $x_{j}^{*}=x_{j}$ and $x_{c}^{*}=x_{c}$ in (1). Another strategy is to use the posterior samples of each child’s best descendant—intuitively, it performs a “lookahead”. In this case, $x_{j}^{*}$ is defined as (similarly for $x_{c}^{*}$):
@@
$x_{j}^{*}=\begin{cases}x_{j}&\text{if }x_{j}\text{ is a leaf node,}\\ \operatorname*{arg\,max}_{x_{c}\in\mathrm{children}(x_{j})}a_{j}(x_{c})&\text{otherwise}.\end{cases}$ (2)
@@
Note that they have the same costs since the posterior samples for both strategies are readily available due to the backup process below, i.e., we do not perform the recursion (2) in this step.
@@
Expansion Given an unexpanded node $x_{i}$, we query the LLM to obtain the top-$k$ most likely children and their corresponding likelihoods. These likelihoods are new observations and we combine them with the prior samples to obtain $n$ posterior samples $(v_{x_{c}})_{n}$ of each child $x_{c}$.
@@
Backup We recursively propagate the newly obtained posterior samples $\{(v_{x_{c}})_{n}\}_{c}$ back up the tree until the root via the path selected in the previous steps. This is done to update the samples contained in each node of the path. The updated nodes’ samples will then influence the selection process in the next iteration, updating the exploration-exploitation tradeoff. Different update strategies can be used depending on the choice of the acquisition function (1). When posterior samples of the children node $x_{c}$ are used to compute $a_{i}(x_{c})$, then we propagate up the posterior samples of the newly expanded nodes as in when computing the prior, i.e. by recursively taking their maximum (Alg. 1). Meanwhile, when the posterior samples of the best descendant of $x_{c}$ are used in $a_{i}(x_{c})$, we simply propagate up the posterior samples of the best child among the newly expanded nodes, without taking further maximums along the path.
@@
Termination criterion  Finally, the posterior samples over the optimal values can not only be used for the selection of new nodes but also to monitor the progress of the optimization. For instance, one can compute the following empirical probability ^$\hat{\mathbb{P}}(c^{*}<v_{\boldsymbol{x}_{0}^{*}})=\frac{1}{N}\sum_{n=1}^{N}\mathbb{I}\left[c^{*}\leq(v_{\boldsymbol{x}_{0}^{*}})_{n}\right]$, which corresponds to the probability of the current best likelihood $c^{*}$ among all leaves ULTS has visited so far is lower than the best value according to our belief $v_{\boldsymbol{x}_{0}^{*}}$. Note that the computation of such a probability is done as in the acquisition function $a(x)$, i.e., using the posterior samples of the root node $\boldsymbol{x}_{0}$ or the posterior samples of the best descendant. Then, one can decide to stop the search once this probability is below some confidence level $\varepsilon>0$.
As a final note, while new nodes are selected greedily, they are selected based on the beliefs over the non-myopic optimal values $v_{x_{i}}$, which take the posterior belief over the likelihood in the remaining steps into account. Moreover, exploration-exploitation is implied since $a_{i}(x_{c})$ is constructed from the posterior samples.
3.5 Practical considerations
@@
In order to put a tractable upper bound on the runtime of ULTS, we introduce a hyperparameter $k_{\max}$ on the maximum number of nodes that can be expanded per level, similar to beam search or top-$k$ sampling. We stop the search as soon as a set of $k_{\max}$ leaves is attained or the termination probability exceeds below $1-\varepsilon$. We shall see in Section 5 that even under this further approximation, we still obtain good results while being efficient in terms of node expansions.
3.6 Limitations
There is reasons to assume that the iid. assumption might not fully hold in practical settings. E.g. the categorical distributions along low-likelihood paths in the search tree might at some point turn into almost uniform distributions (i.e. large $\alpha$) because predicting the next token becomes more and more difficult if many unlikely gibberish tokens are appended to the input prompt. With the restriction of keeping the search computationally tractable, we do not see a straightforward way of gradually varying $\alpha$ across the entire search tree.
In practice, we deal with this by fitting $\alpha$ based on the output along paths with high log-likelihood. This choice is motivated by the following two considerations: First, tokens with very low log-likelihood tend to not be chosen anyways. Miscalibration in low likelihood regions of the search space therefore matter less. Second, there are indications that in general overestimation might be less harmful than underestimation. For example, in A*-type algorithms, the heuristic used should indeed be optimistic [10]. A similar situation is also known for BO methods: Maximum value entropy search [38] also uses an approximation for the distributions over the maximum value which tends to overestimate the true value. They argue (for their setting) that upper bounds still lead to optimization strategies with vanishing regret, whereas this may not hold for lower bounds [39]. For a further discussion of the assumption, see Appendix A.
All in all, our modeling assumptions are chosen with practicality in mind. This is akin to how priors in Bayesian neural networks [41] are chosen, which are usually simply just isotropic Gaussians. Even in Gaussian processes, a generic prior/kernel is often assumed by default [2].
@@ 4 Related Work
Probabilistic/Bayesian optimization on trees have been proposed for game trees. Hennig et al. [11] developed a probabilistic tree search algorithm for game trees, e.g. for playing Tic-Tac-Toe and Go. Crucially, the structure of the game tree is different than the LLM search tree. Moreover, they assumed a Gaussian process prior with the Brownian motion kernel in conjunction with the expectation propagation algorithm [27] to model their beliefs. Grosse et al. [8] extended their approach for a more general directed acyclic graph structure. In contrast, we focus on the tree implied by the sequential generation process in LLMs with a Dirichlet or empirical prior.
Our method can be seen as utilizing a best-first search strategy, of which the A* algorithm [10] is the most famous. Moreover, A*-like beam search algorithm, under the name of best-first beam search, has also been proposed for decoding LLMs [24]. Unlike the non-probabilistic A and best-first beam search, we approach optimization-on-tree problems via the lens of decision-making under uncertainty—putting a prior belief about the unknown, update it based on the observations, and make decision based on the posterior belief. Furthermore, best-first beam search [24] has different goals to ours: it is designed to output the same set of leaves (and thus the optimal likelihood value) as the standard beam search. ULTS, meanwhile, focuses on both performance (i.e., attaining higher likelihood than beam search) and efficiency (reducing the number of costly LLM forward passes).
Finally, Monte-Carlo tree search algorithms, which also utilize samples to make decision, have also been proposed for decoding LLMs [22, 23, 9, 4, 45]. Different from our goal, they focused on incorporating external rewards that are only observable at the leaf nodes. For instance Zhang et al. [45] defined the reward to be the unit-test results for code-generating LLMs. Note further that Monte-Carlo tree search does not define a probabilistic model (i.e. a prior and posterior beliefs) over quantities in the search tree—its decision making is based on the statistics obtained by exploring the search tree on the fly. In contrast, ULTS does not require costly gathering of those statistics during the decoding process and instead makes decisions based on precomputed samples from its beliefs.
@@ 5 Experiments
5.1 Toy Example
As a first experiment, we compare ULTS to beam search on artificially generated search problems from Dirichlet priors. The trees have branching factor $b=8$ and depth $d=5$. Since these trees are so small we optimized the acquisition function in eq. (1) over the the full boundary, i.e. non-recursively. The transition probabilities at each node are sampled from a Dirichlet prior with fixed $\alpha\in\{0.1,0.2,0.5,0.8\}$. The comparison is on-model, i.e., ULTS is run with the ground truth parameter of $\alpha$. We repeat the experiment with different values for the confidence parameter $\varepsilon$ of ULTS from $\{0.05,0.1,0.3\}$. Since the toy problems are so small, the exploration of too many nodes is not an issue and we use $k_{\max}=\infty$. Beam search is run with beam sizes ranging from $1$ to $7$. The results in Fig. 3 show for $\alpha=0.1$ and $\alpha=0.2$ that ULTS dominates across the entire range of hyper-parameters. The results for $\alpha=0.5$ and $\alpha=0.8$ (see Appendix C) show the same pattern. This suggests that knowledge about concentration strength helps reduce the number of search steps.
Figure 3: Comparison on trees, where the transition probabilities are sampled from a Dirichlet prior for two different values of the concentration parameter.
5.2 Experiments with LLMs
We continue with off-model experiments, where we test ULTS for the decoding process of LLMs. We use GPT-2 [29] and Llama-2-7b [33] for text generation on articles from the Wikipedia [31], CNN Daily Mail [13], and Reddit TL;DR [36] datasets. [^2] See Appendix B for URLs to the models and datasets. Unless specified otherwise, we use the strategy in (2) for the selection and backup steps. See Appendix C for results with the other strategy and further dataset/setting. All experiments are done on single NVIDIA GeForce RTX 2080 Ti and NVIDIA A40 48GB GPUs for GPT-2 and Llama-2-7b, respectively.
@@ 5.2 Experiments with LLMs
Since many of the text samples in the Wikipedia dataset end with e.g., references instead of full sentences, we filter for text samples with at least $500$ tokens, resulting in a test set with $332$ token sequences (out of a random subset of originally $1000$ sequences). We use $200$ tokens as input and predict $20$ tokens. We do the same for the CNN Daily Mail dataset, where we end up with $790$ token sequences. We use context length of $300$ and generate $60$ tokens. We also include a summarization task, where the goal is to generate a $40$ token long summary of the input sequence. For this, we use $1000$ random samples from the TL;DR dataset. We use the full context of variable length. Note that we decode sequences of fixed length instead of stopping at the <EOS> token. This makes it easier to compare the different decoding methods with each other, as the overall likelihood of a sequence of course depends on its length. Additionally, see Appendix C for a machine translation experiment where ULTS stops at the <EOS> token.
@@
5.2.1 Choice of the prior
Before the decoding runs, we sampled a set $\mathcal{C}$ of Categorical distributions from the LLM on training sequences for each of the datasets and both of the LLLs. We used $1000$ samples for each of the datasets and LLMs we considered. They are used for the empirical prior, as well as training data for fitting the concentration parameter $\alpha$ of the Dirichlet prior. It is possible to adjust $\alpha$ using maximum likelihood optimization [26]. However, we have found that this did not work very well in our experiments: We suspect that it is either due to numerical problems of the optimization algorithm, or that the best fit for the Dirichlet distribution does not necessarily coincide with the best fit for the implied distribution of the (recursive) maximum. Figure 4 shows samples for the maximum of Categorical distributions returned by the LLM, as well as the distribution of the maximum of the Categorical distribution from a Dirichlet prior for $\alpha=10^{-1},10^{-4},5\times 10^{-6}$. While $\alpha=10^{-1}$ is in the order of the maximum likelihood estimator, the other values seem to fit better.
5.2.2 Comparison with baseline decoding methods
As the baseline we use the commonly-used beam search with beam width $k\in\{1,2,3,4,5,10,20\}$ [32]. These values capture the frequently-used beam widths [40]. Based on the results from the previous section, we run ULTS with both the Dirichlet prior with $\alpha=10^{-4}$ and the empirical prior. Moreover, we set $k_{\max}\in\{2,3,4,5,10,20\}$. We set $\epsilon$ to a default value of $0.1$. Figure 5 shows the results for GPT-2 and Llama-2-7b. The error bars indicate ±$\pm 1$ standard error of the mean based on all test sentences.
@@ 5.2.2 Comparison with baseline decoding methods
No matter the choice of $k_{\max}$ (and $k$), ULTS yields sequences with the same or a higher log-likelihood while expanding fewer nodes. For example in the summarization task with GPT-2, ULTS with the empirical prior and $k_{\text{max}}=20$ achieves an average log-likelihood of $-18.31$ while expanding only $137.90$ nodes. In contrast, beam search with $k=5$ returns sequences of average log-likelihood $-21.33$ even though expanding more nodes (196 nodes). This is the case for both choices of priors, with the empirical prior encouraging exploration a bit more. Figure 3 suggests that the distribution $\max_{i}c_{ji}$ is bimodal and not fully captured by the Dirichlet prior. In particular, our choice of Dirichlet prior is slightly too pessimistic. As a result, the search may stop too early and the available budget may not be fully utilized. However, the budget that is used, is used efficiently.
@@
Our recommendation is thus as follows. When efficiency is the main goal, the Dirichlet prior is preferable—it performs similarly to beam search with smaller beam widths, while being more efficient. However, one must tune the value of $\alpha$. If the search performance is important and an additional tree exploration can be afforded (still more efficient than beam search), then the empirical prior is the best choice—it is also hyperparameter-free.
5.2.3 Runtime

### v2

@@ 3.3 Posterior beliefs over optimal values
Notice that whenever a new node $x_{i}$ is added to the search tree, the likelihood associated with the path from the root to that node $c_{\boldsymbol{x}_{0}\rightarrow x_{i}}$ is fully observed—its distribution is simply a Dirac delta. This means, the joint distribution becomes $p(c_{\boldsymbol{x}_{0}\to x_{i}},\Delta_{i})=\delta(c_{\boldsymbol{x}_{0}\to x_{i}})\,p(\Delta_{i})$. Therefore, to sample $v_{x_{i}}$ given that we have observed $c_{\boldsymbol{x}_{0}\to x_{i}}$—i.e., sampling from the posterior $p(v_{x_{i}}\mid c_{\boldsymbol{x}_{0}\rightarrow x_{i}})$—it is sufficient to sample from $p(\Delta_{i})$ and then simply scale all samples by $c_{\boldsymbol{x}_{0}\rightarrow x_{i}}$. We leverage these posterior samples to make decisions in exploring the tree. Notice that these samples are cheap to get, unlike empirical samples based on simulations/rollouts.
@@ 3.4 Decision making
while ^$\hat{\mathbb{P}}(c^{*}<v_{\boldsymbol{x}_{0}})>\varepsilon$ do
@@
^$\hat{\mathbb{P}}(c^{*}<v_{\boldsymbol{x}_{0}})\leftarrow\frac{1}{N}\sum_{n=1}^{N}\mathbb{I}\left[c^{*}\leq(v_{\boldsymbol{x}_{0}})_{n}\right]$
@@
At each iteration, we use the posterior samples by following the steps of Monte Carlo tree search, but without the expensive rollout step. Alg. 2 contains the pseudocode for the decoding portion of ULTS.
@@
1. Selection Starting from the root node $\boldsymbol{x}_{0}$, we recursively pick a child that has the highest empirical probability to be the best child, as computed based on the posterior samples. I.e. at a node $x_{i}$, we recursively maximize the following acquisition function $a_{i}(x_{c})$ over its children:
@@
$a_{i}(x_{c})=\frac{1}{N}{\displaystyle\sum_{n=1}^{N}}\mathbb{I}\bigl{[}(v_{\hat{x}_{c}})_{n}=\underset{x_{j}\in\mathrm{children}(x_{i})}{\max}(v_{\hat{x}_{j}})_{n}\bigr{]},$
@@
Different realizations of $\hat{x}_{c}$ and $\hat{x}_{j}$ can be used: The most straightforward is to use the children’s posterior samples themselves. This corresponds to simply setting $\hat{x}_{j}=x_{j}$ and $\hat{x}_{c}=x_{c}$. Intuitively, this means that we are estimating the probability that a child $x_{c}$ is the best child among other children of $x_{i}$. Another strategy is to use the posterior samples of each child’s best descendant—intuitively, it performs a “lookahead”. In this case, $\hat{x}_{j}$ is defined as (similarly for $\hat{x}_{c}$):
@@
$\hat{x}_{j}=\begin{cases}x_{j}&\text{if }x_{j}\text{ is a leaf,}\\ \underset{x_{c}\in\mathrm{children}(x_{j})}{\operatorname*{arg\,max}}a_{j}(x_{c})&\text{otherwise}.\end{cases}$
@@
Note that they have the same costs since the posterior samples for both strategies are readily available due to the backup step below, i.e., we do not actually perform the recursion (2) in this step. Furthermore, the acquisition function (1) can be extended further, e.g. to take into account a penalty term (Section 5.4). In any case, this selection process is done until an unexpanded node is selected.
@@
2. Expansion Given an unexpanded node $x_{i}$, we query the LLM to obtain the top-$k$ most likely children and their corresponding likelihoods. These likelihoods are new observations and we combine them with the prior samples to obtain $n$ posterior samples $(v_{x_{c}})_{n}$ of each child $x_{c}$. This step is the most expensive part of any tree-search algorithm since querying the LLM is costly. Akin to BO, ULTS strives to reduce the number of node expansions and thus minimize the number of LLM calls.
@@
3. Backup We recursively propagate the newly obtained posterior samples $\{(v_{x_{c}})_{n}\}_{c}$ back up the tree until the root via the path selected in the previous steps. This is done to update the posterior samples contained in each node of the path. They will then influence the selection process in the next iteration, updating the exploration-exploitation tradeoff. Different update strategies can be used depending on the choice of the acquisition function (1). When posterior samples of the children node $x_{c}$ are used to compute $a_{i}(x_{c})$, then we propagate up the posterior samples of the newly expanded nodes as in when computing the prior, i.e. by recursively taking their maximum. When the posterior samples of the best descendant of $x_{c}$ are used in $a_{i}(x_{c})$, we simply propagate up the posterior samples of the best child among the newly expanded nodes, without taking further maximums along the path.
@@
Termination criterion  Finally, the posterior samples over the optimal values can not only be used for the selection of new nodes but also to monitor the progress of the optimization. For instance, one can compute the following empirical probability ^$\hat{\mathbb{P}}(c^{*}<v_{\boldsymbol{x}_{0}})=\frac{1}{N}\sum_{n=1}^{N}\mathbb{I}\left[c^{*}\leq(v_{\boldsymbol{x}_{0}})_{n}\right]$, which corresponds to the probability of the current best likelihood $c^{*}$ among all leaves ULTS has visited so far is lower than the best value $v_{\boldsymbol{x}_{0}}$ at the root, according to our posterior belief. Note that the computation of such a probability is done as in the acquisition function $a(x)$, i.e., using the posterior samples of the root node $\boldsymbol{x}_{0}$ or the posterior samples of the best descendant. Then, one can decide to stop the search once this probability is below some confidence level $\varepsilon>0$.
3.5 Remarks
Practical considerations  In order to put a tractable upper bound on the runtime of ULTS, we introduce a hyperparameter $k_{\max}$ on the maximum number of nodes that can be expanded per level, similar to beam search or top-$k$ sampling.
@@ 3.5 Remarks
Moreover, we stop the search as soon as a set of $k_{\max}$ leaves is attained or the termination probability exceeds below $1-\varepsilon$. We shall see in Section 5 that even under these further constraints, we still obtain good results while being efficient.
Limitations   There are strong reasons to assume that the iid. assumption might not fully hold.
Indeed, our modeling assumptions are chosen with practicality in mind. This is akin to how priors in Bayesian neural networks (Wilson & Izmailov, 2020) are chosen, which are usually simply just isotropic Gaussians. Even in standard BO, a generic Gaussian-process prior/kernel is often assumed by default (Balandat et al., 2020). For a further discussion of the prior assumptions, see Appendix A. Moreover, our work does not take the miscalibration of LLM softmax outputs into account. Indeed, it has been shown that higher likelihood is not always correlated with human preferences (Holtzman et al., 2019; Wiher et al., 2022). However, LLM calibration is orthogonal to the focus of the present paper—we leave this for future work. Nevertheless, ULTS is extensible: One can construct a different prior or use a different utility function to address this; see Section 5.4 for example.
@@ 4 Related Work
Probabilistic/Bayesian optimization on trees has been proposed for game trees. Hennig et al. (2010) developed a roll-out based probabilistic tree search algorithm for game trees, e.g. for playing Tic-Tac-Toe and Go. Crucially, the structure of the game tree is different than the LLM search tree. Moreover, they assumed a Gaussian process prior with the Brownian motion kernel in conjunction with the expectation propagation algorithm (Minka, 2001) to model their beliefs. Grosse et al. (2021) extended their approach for a more general directed acyclic graph structure. In contrast, we focus on the tree implied by the sequential generation process in LLMs with a Dirichlet or empirical prior.
Our method can be seen as utilizing a best-first search strategy, of which the A* algorithm (Hart et al., 1968) is the most famous. Moreover, an A*-like beam search algorithm, under the name of best-first beam search, has also been proposed for decoding LLMs (Meister et al., 2020). Unlike the non-probabilistic A and best-first beam search, we approach optimization-on-tree problems via the lens of decision-making under uncertainty—putting a prior belief about the unknown, updating it based on the observations, and making decision based on the posterior belief. Furthermore, best-first beam search (Meister et al., 2020) has different goals to ours: it is designed to output the same set of leaves (and thus the optimal likelihood value) as the standard beam search. ULTS, meanwhile, focuses on both performance (i.e., attaining higher likelihood than beam search) and efficiency (reducing the number of costly LLM forward passes).
Finally, Monte Carlo tree search, which also utilizes samples to make decisions, has also been proposed for decoding LLMs (Leblond et al., 2021; Liu et al., 2023; Hao et al., 2023; Feng et al., 2023; Zhang et al., 2023; Zhou et al., 2024; DeLorenzo et al., 2024). Different from our goal, they focused on incorporating external rewards that are only observable at the leaf nodes. For instance, Zhang et al. (2023) defined the reward to be the unit-test results for code-generating LLMs. Moreover, Monte-Carlo tree search does not define a probabilistic model (i.e. a prior and posterior beliefs) over quantities in the search tree—its decision making is based on the statistics obtained by exploring the search tree on the fly. In contrast, ULTS does not require costly gathering of those statistics during the decoding process; instead, it makes decisions based on precomputed samples from its beliefs.
@@ 5 Experiments
Setup  In this section, we evaluate ULTS in both close-ended and open-ended text generation problems. We set ULTS’ $k_{\max}\in\{2,3,4,5,10,20\}$ and set $\epsilon$ to a default value of $0.1$. The error bars indicate ±$\pm 1$ standard error of the mean based on all test sentences. We compare ULTS against beam search (the de-facto decoding method for machine translation), as well as the recent baselines multinomial beam search (Multinomial BS; Kool et al., 2019), contrastive search Su et al. (2022) nucleus sampling (Holtzman et al., 2019), best-of-k sampling (Stiennon et al., 2020), speculative decoding (Leviathan et al., 2023), and DoLA (Chuang et al., 2024). For beam search, multinomial beam search, and best-of-k baselines, we evaluate beam sizes and numbers of samples $k\in\{1,2,3,4,5,10,20\}$, respectively. We set any other hyperparameters of all baselines as suggested by the original papers or by Huggingface. We use Huggingface’s implementation of the baselines and provide our implementation of ULTS in the supplements.
Choice of the prior  We sampled a set $\mathcal{C}$ of Categorical distributions from the LLM on training sequences for each of the datasets and both of the LLMs. We used $1000$ samples from the training set for each of the datasets and LLMs we considered. They are used for the empirical prior, as well as training data for fitting the concentration parameter $\alpha$ of the Dirichlet prior.
Figure 7 in Section C.3 shows samples for the maximum of Categorical distributions returned by the LLM, as well as the distribution of the maximum of the Categorical distribution from a Dirichlet prior for $\alpha=10^{-1},10^{-4},5\times 10^{-6}$. For the experiments below, we picked a value for $\alpha$ based on the histograms in Figure 7. For a comparison between different values of the concentration parameter, please refer to Fig. 8 (Section C.3). Note that this way of choosing the hyperparameter $\alpha$ is very convenient compared to performing costly cross-validation.
Figure 3: Left: Machine translation results with the WMT-19 English-to-German dataset in term of log-likelihood. Right: The corresponding BLEU scores.
5.1 Close-ended generation
We compare ULTS with the baselines in a machine translation task with 1000 randomly sampled sequences from the WMT-19 English to German dataset in Fig. 3 in terms of likelihood and BLEU score. For the LLM, we use T5-large (Raffel et al., 2020); for ULTS, we use a Dirichlet prior with $\alpha=5\times 10^{6}$. Unless specified otherwise, we use the strategy in (2) for the selection and backup steps. See Appendix C for results with the other strategy and further dataset/setting. All experiments are done on single NVIDIA GeForce RTX 2080 Ti and NVIDIA A40 48GB GPUs for GPT-2 and Llama-2-7b, respectively. We stop the exploration of a path in the tree the <EOS> token is found or when the maximum depth of 60 is exceeded.
The results are in Figs. 3 and 3. ULTS are both more efficient (fewer node expansion) and more performant (higher average log-likelihood) at all values of $k$/$k_{\text{max}}$. For all methods, small to medium beam sizes of $k=3$ seem to work best in terms of BLEU scores—these scores decrease for a larger (maximum) beam with, in contrast to the log-likelihood in Fig. 3 . See Section C.2, for qualitative examples of the translated sentences.
5.2 Open-ended generation
@@
We use GPT-2 (Radford et al., 2019) and Llama-2-7b (Touvron et al., 2023) for text generation on articles from the Wikipedia (See et al., 2017), CNN Daily Mail (Hermann et al., 2015), and Reddit TL;DR (Völske et al., 2017) datasets. Since many of the text samples in the Wikipedia dataset end with e.g., references instead of full sentences, we filter for text samples with at least $500$ tokens, resulting in a test set with $332$ token sequences (out of a random subset of originally $1000$ sequences). We use $200$ tokens as input and predict $20$ tokens. We do the same for the CNN Daily Mail dataset, where we end up with $790$ token sequences. We use context length of $300$ and generate $60$ tokens. We also include a summarization task, where the goal is to generate a $40$ token long summary of the input sequence. For this, we use $1000$ random samples from the TL;DR dataset with variable-length contexts. Here, we decode sequences of fixed length instead of stopping at the <EOS> token. We run ULTS with both the Dirichlet prior with $\alpha=10^{-4}$ and the empirical prior.
@@
Figure 4 shows the results. No matter the choice of $k_{\max}$ (and $k$), ULTS yields sequences with the same or a higher log-likelihood while expanding fewer nodes. For example in the summarization task with GPT-2, ULTS with the empirical prior and $k_{\text{max}}=20$ achieves an average log-likelihood of $-18.31$ while expanding only $137.90$ nodes. In contrast, beam search with $k=5$ returns sequences of average log-likelihood $-21.33$ even though expanding more nodes (196 nodes). This is the case for both choices of priors, with the empirical prior encouraging exploration a bit more. The histogram in Fig. 6 (Section C.3) suggests that the distribution $\max_{i}c_{ji}$ is bimodal and not fully captured by the Dirichlet prior. In particular, our choice of Dirichlet prior is slightly too pessimistic. As a result, the search may stop too early and the available budget may not be fully utilized. However, the budget that is used, is used efficiently.
@@
Our recommendation is thus as follows. When efficiency is the main goal, a Dirichlet prior with low concentration is preferable—it performs similarly to beam search with smaller beam widths, while being more efficient. If the search performance is important and an additional tree exploration can be afforded (still more efficient than beam search), then the empirical prior is the best choice—it is also hyperparameter-free.
5.3 Runtime